A bowling ball that has an 11-cm radius and a 7.2-kg mass is rolling without slipping at 2.0 m/s on a horizontal ball return. It continues to roll without slipping up a hill to a height h before momentarily coming to rest and then rolling back down the hill. Model the bowling ball as a uniform sphere and find h.
1
Expert's answer
2020-11-25T10:53:09-0500

According to the law of energy conservation kinetic energy of the rolling ball is equal to potential energy of the ball on height h:

"E_{k}=E_{p}"

"\frac{m\times V^2}{2}+\frac{I\times \omega^2}{2}=m\times g\times h"

I - moment of inertia of the ball

"I=\frac{2}{5}\times m \times r^2"

where m - mass, r - radius.

"\omega" is anglular velocity

"\omega = \frac{V}{r}"

calculating

"\omega = \frac{2}{0.11}=18.18"

"I=\frac{2}{5}\times 7.2 \times 0.11^2=0.035"

"\frac{m\times V^2+I\times \omega^2}{2}=m\times g\times h"

"h=\frac{m\times V^2+I\times \omega^2}{2\times m \times g}"

"h=\frac{7.2\times 2^2+0.035\times 18.18^2}{2\times 7.2 \times 9.8}=\frac{28.8+11.57}{141.12}=0.29m"